OpenStudy (anonymous):

Two bus drivers , A and B , return to bus terminal every 50 and 140 minutes respectively . If they leave the bus terminal at 0630 , when will they next meet at bus terminal . I understand it but i got 1710 h but the answer sheet wrote1100h . Please helppp!!

OpenStudy (anonymous):

This is a least common multiple problem Find the LCM of 50 and 140 and that will be your answer. Do you need assistance finding the LCM of 50 and 140 ?

OpenStudy (anonymous):

Rather that will be the number of minutes after 0630 that they will meet in the terminal..Then you just add that time to 0630 to get your answer.

OpenStudy (anonymous):

the least common multiple is 700 . isn't it?

OpenStudy (anonymous):

50=2x5^2 140=2^2x5x7 2^2x5^2x7=700

OpenStudy (anonymous):

yeah... 700 = 60*11 + 40 so that's 11 hours and 40 min 0630 + 11 hours 40 min = 1730 + 40 min = 1810

OpenStudy (anonymous):

oh yah . thank you . So the answer will be 1810 and not 1100 right ?

OpenStudy (anonymous):

yes.